OpenStudy (anonymous):

Write an expression for the area of the rectangle in terms of x. Area 63, length x+2 width x

jhonyy9 (jhonyy9):

do you know formula for area of a rectangle ?

OpenStudy (anonymous):

yep length times width

jhonyy9 (jhonyy9):

yes so than will be area = length time width so A =(x+2)*x so 63=(x+2)*x can you continue ?

jhonyy9 (jhonyy9):

x^2 +2x =63 x^2 +2x -63 = 0 can you factoriz it completly ?

jhonyy9 (jhonyy9):

x^2 +2x -63 = 0 (x+9)(x-7) =0 can you solve it now ?

OpenStudy (anonymous):

yh thamk you for that

OpenStudy (anonymous):

*thank

jhonyy9 (jhonyy9):

x+9=0 x=-9 this mean that the length equal 9 x-7=0 x=7 this result the width equal 7

jhonyy9 (jhonyy9):

because 63=9*7

jhonyy9 (jhonyy9):

do you understand it now sure ?